OpenStudy (anonymous):

Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (3, 6) and (–2, –4). (4 points) Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

OpenStudy (anonymous):

Line equation through two points A(x1,y1) and B(x2,y2) is given by y-y1=(y2-y1)/(x2-x1) (x-x1) in your case y-6=(-4-6)/(-2-3) (x-3) y-6=2(x-3) y=6+2x-6 y=2x slope–intercept form. is better cause u see slope and intersept point and in standard form all is mixed up

OpenStudy (anonymous):

that was Part 1